Find the dimensions of a right circular cone of minimum volume V that can be

circumscribed about a sphere of radius 8 inches.


1
Expert's answer
2022-03-15T17:12:19-0400




The volume of a cone = "\u200b\n \\tfrac{1}{3}*\u03c0 *r^2*h"

Let the radius of the sphere = a

the radius of the cone =r

by similar triangles "sin (θ)=\tfrac{a}{h-a}" = "\\tfrac{r}{\\sqrt{h^2+r^2}}\n\u200b\n \n\u200b\n \n\u200b"


"=a*(\sqrt{h^2+r^2) }=r(h-a)"

on squaring both sides

"a^2(h^2+r^2)=r^2(h-a)^2"

"=a^2h^2+a^2r^2= h^2r^2-2ahr^2+a^2r^2"

making "r^2" to be the subject of the formulae we have

"r^2= \frac{a^2h}{h-2a}"

The volume V = "\u200b\n \\tfrac{1}{3}*\u03c0 *\\frac{h^2}{h-2a}*h"

at minimum volume ;


"\frac{dV}{dh}= \frac{\pi a^2}{3}[\frac{(h-2a)2h-h^2(1)}{(h-2a)^2}]=0"

"2(h-2a)h-h^2=0"


"2h-4a-h=0"

"h-4a=0"

"h=4a"

but a =8 inches

therefore "h=4*8=32"

The height of the cone is 32 inches.

"r^2= \frac{8^2*32}{32-2*8}"


"r^2=\frac{2048}{16}"


"r^2=128"


"r= \sqrt{128}"


"r= 11.3137"


The height of the cone is 32 inches and the radius is 11.3137inches.